# An Empirical Study of Catastrophic Forgetting in Large Language Models During Continual Fine-tuning

### v4

@@ Abstract
Catastrophic forgetting (CF) is a phenomenon that occurs in machine learning when a model forgets previously learned information while acquiring new knowledge for achieving a satisfactory performance in downstream tasks. As large language models (LLMs) have demonstrated remarkable performance, it is intriguing to investigate whether CF exists during the continual instruction tuning of LLMs. This study empirically evaluates the forgetting phenomenon in LLMs’ knowledge during continual instruction tuning from the perspectives of domain knowledge, reasoning, and reading comprehension. The experiments reveal that catastrophic forgetting is generally observed in LLMs ranging from 1b to 7b parameters. Surprisingly, as the model scale increases, the severity of forgetting intensifies in such a model sale range which may result from the much significant initial performance in the larger LLM. Comparing the decoder-only model BLOOMZ with the encoder-decoder model mT0, BLOOMZ exhibits less forgetting and retains more knowledge. Interestingly, we also observe that LLMs can mitigate language biases, such as gender bias, during continual fine-tuning. Furthermore, our findings indicate that general instruction tuning can help alleviate the forgetting phenomenon in LLMs during subsequent fine-tuning.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive general capabilities in tackling a wide range of tasks (Bubeck et al., 2023). However, when it comes to real-world applications, users often find that certain specific abilities need enhancement. In such cases, relevant task-specific data are adopted to fine-tune the model in instructions to improve its performance in particular tasks (Touvron et al., 2023; Scialom et al., 2022). However, the widely used LLMs such as LLAMA-Chat (Touvron et al., 2023), ChatGPT [^1] https://openai.com/, and Claude-2  [^2] https://www.anthropic.com/news/claude-2 have been fine-tuned, and the continual instruction tuning in specific tasks may affect the general knowledge stored in LLMs since the original training data are inaccessible. The study of continual instruction tuning is crucial for the practical application of LLMs. For example, in Table 1, an LLM is first fine-tuned for the task of title generation and then learns to generate an explanation. This scenario is related to the continual learning paradigm in machine learning, where a model is trained on a sequence of tasks (Lopez-Paz & Ranzato, 2017; Wu et al., 2022; Ke & Liu, 2022). A significant challenge in this paradigm is catastrophic forgetting (CF), in which a model forgets previously learned knowledge due to parameter updates during the learning of new tasks, leading to decreased performance on earlier tasks (Li & Hoiem, 2017; Lopez-Paz & Ranzato, 2017).

### v5

@@ Abstract
Catastrophic forgetting (CF) is a phenomenon that occurs in machine learning when a model forgets previously learned information while acquiring new knowledge to achieve satisfactory performance in downstream tasks. As large language models (LLMs) have demonstrated remarkable performance, it is intriguing to investigate whether CF exists during the continual instruction tuning of LLMs. This study empirically evaluates the forgetting phenomenon in LLMs’ knowledge during continual instruction tuning from the perspectives of domain knowledge, reasoning, and reading comprehension. The experiments reveal that catastrophic forgetting is generally observed in LLMs ranging from 1b to 7b parameters. Surprisingly, as the model scale increases, the severity of forgetting intensifies in such a model sale range which may result from the much significant initial performance in the larger LLM. When comparing the BLOOMZ decoder-only model with the encoder-decoder model mT0, BLOOMZ exhibits less forgetting and retains more knowledge. Interestingly, we also observe that LLMs can mitigate language biases, such as gender bias, during continual fine-tuning. Furthermore, our findings indicate that general instruction tuning can help alleviate the forgetting phenomenon in LLMs during subsequent fine-tuning.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive general capabilities in tackling a wide range of tasks (Bubeck et al., 2023). However, when it comes to real-world applications, users often find that certain specific abilities need enhancement. In such cases, relevant task-specific data are adopted to fine-tune the model in instructions to improve its performance in particular tasks (Touvron et al., 2023; Scialom et al., 2022). However, the widely used LLMs such as LLAMA-Chat (Touvron et al., 2023), ChatGPT [^1] https://openai.com/, and Claude-2  [^2] https://www.anthropic.com/news/claude-2 have been fine-tuned, and continual instruction tuning in specific tasks may affect the general knowledge stored in LLMs since the original training data are inaccessible. The study of continual instruction tuning is crucial for the practical application of LLMs. For example, in Table 1, an LLM is first fine-tuned for the task of title generation and then learns to generate an explanation. This scenario is related to the continual learning paradigm in machine learning, where a model is trained on a sequence of tasks (Lopez-Paz & Ranzato, 2017; Wu et al., 2022; Ke & Liu, 2022). A significant challenge in this paradigm is catastrophic forgetting (CF), in which a model forgets previously learned knowledge due to parameter updates during the learning of new tasks, leading to decreased performance on earlier tasks (Li & Hoiem, 2017; Lopez-Paz & Ranzato, 2017).